# EquiTabPFN: A Target-Permutation Equivariant Prior Fitted Networks

### v1

@@ 4.1 Target permutation equivariance
When presented with a new unseen dataset of covariate/target pairs $(x_{n},y_{n})_{n=1}^{N}$, the order of the component’s target is arbitrary. In other words, given target vectors of the form $y_{n}=\left((y_{n})_{1},\dots,(y_{n})_{q}\right)$, these could as well be presented in a different order by applying a permutation $\sigma$ to the components of $y_{n}$ to obtain a permuted target $\sigma(y_{n})=\left((y_{n})_{\sigma(1)},\dots,(y_{n})_{\sigma(q)}\right)$. The transformed dataset $(x_{n},\sigma(y_{n}))_{n=1}^{N}$ is still essentially the same as the original dataset up to the permutation as we only changed the order of the target components. For instance, when the target represents a one hot encoding vector of $2$ classes: ”red” or ”blue”, it should not matter whether we encode ”red” as the first or the second class in the one-hot vector. Consequently, a pre-trained model should be able to provide consistent predictions regardless of the component’s order. More formally, the model should satisfy the following equivariance property:
@@
4.2 Non-equivariance of TabPFN model
As introduced in Hollmann et al. (2023), TabPFN is not permutation equivariant in the target’s components. Consequently, it is not guaranteed to provide consistent predictions when the target components are permuted, thus affecting its robustness. To illustrate the implications of non-equivariance on the robustness of TabPFN, we consider a toy classification problem in $2$ dimensions, where $9$ training covariates are positioned on a $2$-dimensional regular grid, each corresponding to a different class inspired by McCarter (2024). The pre-trained TabPFN model is then used to predict the classes on a regular grid of 900 points. Figure 1 (top) shows the classification boundaries when using the same data but with $3$ different orderings for the classes. It is clear from the figure that the ordering heavily affects the prediction even in this simple example, which strongly impacts robustness of models like TabPFN. While Hollmann et al. (2023) propose a mitigation strategy by averaging several predictions using random permutations of the target dimensions, that approach is costly and can be impractical when the output dimension $q$ is large since $\mathcal{O}(q!)$ forward predictions are required to make the function equivariant as discussed in the next section.
The non-equivariance of TabPFN is mainly due to the choice of both encoders and decoders that force an arbitrary order on the dimensions of the targets. In Section 5, we overcome this limitation by introducing a new architecture that is target permutation equivariant. As illustrated in Figure 1 (bottom), the predictions made by the proposed network remain the same regardless of the class ordering. Next, we provide a theoretical justification for enforcing equivariance in the model by analyzing the optimal solution to the pre-training objective considered in Equation 1.
4.3 Bounds on non-equivariant models
From a theoretical standpoint, permutation equivariance of a pre-trained model is also a natural and important property to satisfy. To illustrate this fact, we start with a decomposition of the objective into an equivariance gap and an optimality error and then establish that the optimal solution of the pre-training problem in Equation 1 must be permutation equivariant in the target’s components, as soon as the model is expressive enough.
Equivariance gap. We define the equivariance gap $\mathcal{E}^{\text{equi}}(f)$ of a function $f$ relative to $\mathcal{L}$ as the difference between the objective value at $f$ and at its symmetrized version ${f}^{\text{equi}}$:
$\mathcal{E}^{\text{equi}}(f):=\mathcal{L}(f)-\mathcal{L}({f}^{\text{equi}}).$ (3)
Here, the symmetrized function ${f}^{\text{equi}}$ is an equivariant function obtained from $f$ by the averaging operation:
${f}^{\text{equi}}_{X,Y}\left(X^{\star}\right)=\mathbb{E}_{\sigma}\left[\sigma^{-1}\left(f_{X,\sigma(Y)}\left(X^{\star}\right)\right)\right],$ (4)
where $\mathbb{E}_{\sigma}$ denotes the expectation w.r.t. to the uniform distribution over all permutations $\sigma$. By construction, ${f}^{\text{equi}}$ is permutation equivariant w.r.t. the target components, meaning that it satisfies Equation 2 in Definition 4.1. Moreover, a function $f$ is itself equivariant iff ${f}^{\text{equi}}=f$. Hence, the gap $\mathcal{E}^{\text{equi}}(f)$ vanishes as soon as $f$ is itself equivariant.
inline]DS: I wonder if we should reintroduce the proof for Equation 4 as it is pedagogic and I suspect some readers will find it useful/struggle to derive it themselves.
In general, the equivariance gap can take negative values. However, we establish later that this equivariance gap must be non-negative under the following assumptions on the pointwise loss $\ell$ and the marginal distribution $p$ of the data:
@@ 4.3 Bounds on non-equivariant models
    Invariance and convexity of the pointwise loss. The pointwise loss $\ell$ is strictly convex in its first argument and is invariant to permutations of the components of its arguments, i.e. for any permutation $\sigma$, it holds that:
    $\ell(\sigma(y),\sigma(y^{\prime}))=\ell(y,y^{\prime}).$
@@
    Invariance of the data distribution. The marginal distribution of the data is invariant to permutations applied to $Y$ and $Y^{\star}$, i.e.
    $p(X,\sigma(Y),X^{\star},\sigma(Y^{\star}))=p(X,Y,X^{\star},Y^{\star}).$
@@
Under Assumptions (A) and (B), the objective $\mathcal{L}(f)$ in Equation 1 admits the decomposition:
$\mathcal{L}(f)=\mathcal{L}({f}^{\text{equi}})+\mathcal{E}^{\text{equi}}(f),$ (5)
where the equivariance gap $\mathcal{E}^{\text{equi}}(f)$ is always non-negative and only equal to 00 when $f$ is equivariant to permutations. Furthermore, when $\ell$ is the squared loss (i.e. $\ell(y,y^{\prime})=\lVert y-y^{\prime}\rVert^{2}$), then
$\mathcal{E}^{\text{equi}}(f)=\mathbb{E}_{p}\left[\lVert f_{X,Y}\left(X^{\star}\right)-{f}^{\text{equi}}_{X,Y}\left(X^{\star}\right)\rVert^{2}\right].$
@@
The key step is to express the objective as an expectation over permutations using the invariance assumptions Assumptions (A) and (B) on the data and loss:
$\mathcal{L}(f)=\mathbb{E}_{p}\mathbb{E}_{\sigma}\left[\ell\left(\sigma^{-1}f_{X,\sigma(Y)}\left(X^{\star}\right),Y^{\star}\right)\right].$
The non-negativity of the equivariance gap is then established using Jensen’s inequality by convexity of the loss $\ell$ in its first argument (Assumption (A)). Finally, the expression of the gap for the quadratic objective follows by explicit calculations. ∎
@@
Proposition 4.2 shows that minimizing the objective $\mathcal{L}$ requires reducing both the equivariance gap and the optimality error of symmetrized functions. Hence, the model must employ some of its expressive power solely for being equivariant, which can be wasteful. This fact is further emphasized by the following proposition which shows that if $\mathcal{F}$ is rich enough, any minimizer of $\mathcal{L}$ over $\mathcal{F}$ must be equivariant.
Proposition 4.3.
Assume that $\mathcal{F}$ is rich enough so that for any $f\in\mathcal{F}$, then ${f}^{\text{equi}}$ also belongs to $\mathcal{F}$. Then, under Assumptions (A) and (B), any minimizer $f^{\star}$ of $\mathcal{L}$ over $\mathcal{F}$ must be equivariant.
Proof.
Let $f^{\star}$ be a minimizer of $\mathcal{L}$ over $\mathcal{F}$. Assume by contradiction that $f^{\star}$ is not equivariant. Hence, by Proposition 4.2, it follows that $\mathcal{E}^{\text{equi}}(f^{\star})>0$. Moreover, the error decomposition $\mathcal{L}(f^{\star})=\mathcal{L}(\tilde{f^{\star}})+\mathcal{E}^{\text{equi}}(f^{\star})$ from Proposition 4.2 ensures that $\mathcal{L}(f^{\star})>\mathcal{L}({f^{\star}}^{\text{equi}})$. However, by assumption on $\mathcal{F}$, the function ${f^{\star}}^{\text{equi}}$ belongs to $\mathcal{F}$, which contradicts the optimality of $f^{\star}$. By contradiction, we conclude that $f^{\star}$ must be equivariant. ∎
A consequence of Proposition 4.3 is that a non-equivariant model must learn to be equivariant in addition to reducing the prediction error, as we further verify empirically in Section 6. Next, we introduce an equivariant architecture that bypasses these limitations while resulting in competitive performance.
5 Target Equivariant Prior Fitted Network
Figure 2: Overview of EquiTabPFN’s architecture. Data is tokenized via an encoder, processed using self-attention, and decoded to obtain predictions. The encoder maps covariates to a single token and embeds target components into tokens via a $1\times 1$ convolution. Missing test tokens are replaced by prediction tokens. Self-attention alternates between two blocks: (1) Feature-wise attention, where target tokens attend only to covariate tokens (gray arrows), and covariate tokens attend to all tokens (blue arrows); (2) Data-wise attention, where test tokens attend only to training tokens (blue arrows), and training tokens attend to themselves (gray arrows). The decoder applies cross-attention between test and training tokens (queries $Q$, keys $K$), using original target vectors as values $V$, followed by a residual connection to preserve equivariance. The final output $f_{X,Y}(X^{\star})$ is compared to the test target $Y^{\star}$ in the training objective.
inline]Caption for Figure 2
We introduce EquiTabPFN, a new model architecture for in-context learning on tabular data, that is permutation equivariant w.r.t. the target’s components. Our architecture integrates self-attention mechanisms across data points and data components to leverage relationships between datapoints while preserving equivariance by processing individual attributes (such as the targets components). Following the notation in Section 3, EquiTabPFN takes training covariate/target pairs $(X,Y)$ and test covariate $X^{\star}$ as inputs and produces a prediction $\hat{Y}=f_{X,Y}\left(X^{\star}\right)$ of the test target $Y^{\star}$. Unlike TabPFN which requires fixed target dimensions for all datasets, EquiTabPFN allows the dimensions of the target to change depending on the dataset as a consequence of its equivariant architecture. EquiTabPFN consists of four major modules: a target equivariant encoder, an attention module both across data components and across data points, and a non-parametric equivariant decoder, each designed to facilitate the end-to-end learning of components interactions and datapoint relationships, see Figure 2. Below, we elaborate on each module and their interactions.
5.1 Target equivariant encoder
The encoder constructs training and test tokens by applying a linear projection to both covariates and targets so that target equivariance is preserved. Specifically, following (Hollmann et al., 2023), each training and test covariate vector $x_{n}$ and $x^{\star}_{m}$ is encoded into a single token of dimension $d$ by applying a linear projection matrix $U$ of size $d\times p$. However, instead of adding a linear projection of the training targets to each corresponding training token, as done in the case of TabPFN (see Section 3), we compute a token for each component $(y_{n})_{j}$ of a training target by multiplying them with an embedding vector $V$ of dimension $d$ for all $1\leq j\leq q$. This operation amounts to applying a $1\times 1$ convolution along the components of each target which preserves target equivariance. Since, the validation target $Y^{\star}$ is not provided to the model as input, it is replaced by a trainable prediction token $W_{pred}$ of dimension $d$ that is repeated $M\times q$ times to form an initial guess $\tilde{Y}^{0}$ of the target. All these embeddings along with prediction tokens are collected to form a single tensor $E$ of shape $(N+M,q+1,d)$, where the blocks $E_{:N,1,:}$, $E_{N:M,1,:}$ correspond to embeddings of $X$ and $X^{\star}$, $E_{:N,1:q,:}$ represents the embedding of $Y$ while $E_{N:M,1:q,:}$ denotes the initial guess $\tilde{Y}^{0}$ obtained using the prediction token. This tensor is then processed by the attention modules as described next. inline]DS: Write matrix dimensions
5.2 Self-Attention Mechanisms
The core of the architecture involves two alternating self-attention modules: self-attention across components $\text{SelfAtt}_{c}$ and self-attention across datapoints $\text{SelfAtt}_{b}$ used for transforming the tokens. These alternating self-attention layers allow the model to learn both intra-samples components interactions and inter-samples relationships. Following standard design choices for transformers (Vaswani et al., 2017), we apply residual connections and layer normalization to ensure stability and robust gradient flow, i.e.:
$E$ $\leftarrow\text{LN}\left(E+\textbf{SelfAtt}_{c/b}(E)\right),$
$E$ $\leftarrow\text{LN}\left(E+\text{MLP}(E)\right),$
where LN denotes the layer normalization layer (Ba et al., 2016), $\text{SelfAtt}_{c/b}$ denotes one of the considered self-attention mechanisms and MLP is a one hidden-layer network acting on each embedding independently. Below, we describe both self-attention mechanisms in more detail.
Self-attention across target components
Self-Attention across datapoints
captures relationships between datapoint embeddings, allowing the model to aggregate information globally. It is applied between samples and independently per each input dimensions $p$ and $q$ to preserve equivariance. Similarly to (Hollmann et al., 2023), training and validation tokens only attend to training tokens. This is ensured in practice applying a suitable mask when computing attention.
5.3 Non-parametric equivariant decoder
The decoder aggregates the processed embeddings to produce prediction $\hat{Y}$. This is achieved in two steps: an attention module first computes an intermediate prediction $\tilde{Y}=(\tilde{y}_{m})_{m=1}^{M}$ in the form of a weighted average of training targets $Y$, then a residual correction is added to produce the final prediction. More precisely, the attention module uses the embeddings of the training and validation samples as keys and queries, while the attention values are simply the training targets $Y$:
$\tilde{y}_{m}=\sum_{n=1}^{N}y_{n}\text{SoftMax}\left(\frac{\sum_{i,u}E_{n,i,u}E_{m,i,u}}{\sqrt{(1+q)d}}\right).$ (6)
The residual correction, in the form of a point-wise MLP, operates independently on each dimension $j$ of the attention output $\tilde{y}_{m}$ so that equivariance is preserved while enabling nonlinear interactions between training values.
Without the residual correction and removing the dependence of the keys and queries embeddings on the training targets $Y$ (for instance by setting the weights of the target encoder and pointwise MLP to 00), the decoder becomes a linear non-parametric regression estimator (Tsybakov, 2009, Definition 1.7), which is a generalization of Nadaraya-Watson’s estimator  [^1] Linearity, here, is w.r.t. the targets $Y$. (Nadaraya, 1964; Watson, 1964). However, linear estimators are known to be suboptimal compared to non-linear ones (Donoho & Johnstone, 1998). This motivates introducing a nonlinear dependence of the estimator to $Y$, in our setting, to increase the expressiveness of the decoder allowing it to adapt to the prediction task at hand. Experimentally, we also found a clear improvement when adding such a residual correction and making the embeddings dependent on the targets.
@@ 6 Experiments
Figure 3: Equivariance error for TabPFN observed while training (left) and after training when considering different number of classes and ensembles (right). Training reduces slowly the equivariance error but even at the end of the pre-training the model is not close to being equivariant, even when considering ensembles, in particular when considering more classes.
We perform numerical experiments to evaluate our proposed model and compare it with two state-of-the-art pre-trained models for tabular dataset (TabPFN (Hollmann et al., 2023) and MotherNet (Müller et al., 2023)), as well as classical machine learning methods.
@@ 6.1 Experimental setup
Training procedure. We use a similar training protocol as in Hollmann et al. (2023), in which the model is trained on classification datasets generated according to their proposed artificial dataset prior. In this protocol, each dataset has a fixed size of $1024$ and is split into training and test uniformly at random. The maximum number of classes is fixed to $10$, while the maximum dimension of the covariate vector is fixed to $100$. Following Müller et al. (2023), we represent the target $y$ as a one-hot encoding vector whose dimension is the number of classes in the dataset. Moreover, we employ the exact same strategy for handling missing values in the covariates. Training is performed using 153600 batches of 72 synthetically generated datasets each, which means the model was exposed to $\sim$11M artificial datasets during pre-training. This is a similar order of magnitude of datasets used for pre-training TabPFN by Hollmann et al. (2023). We used the Adam optimizer (Kingma & Ba, 2015) with initial learning rate of $0.0001$ and linear-warmup scheduler for the first $10$ epochs followed by cosine annealing (Loshchilov & Hutter, 2017) as in Hollmann et al. (2023). The total training time of the network lasts approximately 4 days on a single A100 GPU with 80GB of GPU memory. The resulting network is then used for all our evaluations without altering its parameters.
Architecture details. We use an EquiTabPFN network with $12$ self-attention layers alternating between both type of attention introduced in Section 5: 6 blocks $\text{SelfAtt}_{c}$ and 6 blocks $\text{SelfAtt}_{b}$. Each self-attention layer consists of a multi-head attention blocks with $4$ heads, embeddings of dimension $512$, and hidden layers of dimension $1024$. This choice ensures a fair comparison with the models used in Hollmann et al. (2023); Müller et al. (2023), since the number of parameters ($25.17M$) are of the same order when counting them as proposed in Kaplan et al. (2020).
6.2 Evaluation of the equivariance gap
We re-trained TabPFN using the same training protocol as Hollmann et al. (2023) to then evaluate its equivariance gap at intermediate training steps. Figure 3 (left) shows the equivariance error of TabPFN in terms of percentage of violation of Equation 2, e.g. how frequently the predicted classes $f_{X,\sigma(Y)}(X^{*})$ and $\sigma\left(f_{X,Y}(X^{*})\right)$ differ. We use over 512 random datasets drawn from the prior and report standard deviation. We do not make those plots for EquiTabPFN as the equivariance error is exactly 0 by construction.
In Figure 3 (left), the equivariance error is clear and slowly decreases during training. This non-equivariance 1) induces additional error for the model as demonstrated in Proposition 4.2 and 2) causes the model to provide surprising predictions given that permuting the output order can change the results as seen in Figure 1.
To avoid this issue, Hollmann et al. (2023) proposed to run multiple predictions while randomly permuting the output. If done with all possible permutations, this gives an equivariant function as discussed in Equation 4 however, this requires making $\mathcal{O}(q!)$ forward passes where $q$ is the number of features. This becomes quickly prohibitive, even for $q=10$ as considered in the original study. We illustrate how fast the model becomes equivariant when using ensembles in Figure 3 (right). This clearly shows that while ensembling more permutation helps to make the model more equivariant, many ensembles are required, in particular when considering more classes as one needs to approximate the expectation of Eq. 4 which has $\mathcal{O}(q!)$ terms.
6.3 Evaluation on real-world datasets
Following Hollmann et al. (2023), we evaluate EquiTabPFN on 30 real-world classification datasets from the OpenML-CC18 benchmark suite (Bischl et al., 2021) that contain less than $2000$ samples, $100$ covariates and $10$ classes (see Table 3 of the appendix for a list of the 30 datasets). Table 1 reports the average AUC on these datasets.
We compare EquiTabPFN with two types of models: recent SOTA pre-trained models and more traditional machine learning methods. For pre-trained models, we use the ones provided by TabPFN (Hollmann et al., 2023) and MotherNet (Müller et al., 2023) as they achieved state-of-the-art results. We refer to these works for an in-depth comparison with other recent AutoML methods. For the traditional baselines, we consider similar baselines as Müller et al. (2023): Histogram Gradient Boosting (XGBoost) (Chen & Guestrin, 2016), Logistic Regression, k-Nearest Neighbors, Random Forests and a vanilla MLP. We refer to that work for the details of the hyper-parameters for these baselines.
The method we propose outperforms TabPFN as well as other baselines which illustrates the practical benefit of having a target-equivariant model. We also report the performance of TabPFN and EquiTabPFN when using ensembling (method + Ens). Ensembling is done by averaging the predictions over several subsamples of the dataset and permutations of the target with the same procedure proposed by Hollmann et al. (2023). Our method also benefits from ensembling but less so than TabPFN since the equivariance of the target is built in the inductive bias of our architecture.
In Figure 4, we show the average rank for each method together with horizontal bars to denote methods that are tied statistically (Demšar, 2006). While our method outperforms TabPFN in average rank and AUC, it is statistically tied with TabPFN but is the only baseline that is not tied statistically with XGBoost.
Model Average AUC
EquiTabPFN $+$ Ens 0.894851
TabPFN $+$ Ens 0.893162
Mothernet $+$ Ens 0.886897
EquiTabPFN 0.893040
TabPFN 0.890675
XGBoost 0.885853
Logistic regression 0.883984
Mothernet 0.879750
Random Forest 0.876427
K-nearest neighbors 0.848294
MLP 0.788029
Table 1: Average AUC over 30 real-world datasets from OpenML-CC18 benchmark suite with less than $10$ classes.
Figure 4: Critical diagram on the 30 real-world datasets from OpenML-CC18 benchmark.
6.4 Evaluation on datasets with unseen class counts
Model Averaged AUC
EquiTabPFN 0.950626
Random Forest 0.942480
XGBoost 0.938711
Logistic regression 0.921125
K-nearest neighbors 0.872002
Table 2: Average AUC over 9 real-world datasets from OpenML-CC18 benchmark suite with more than $10$ classes.
Table 2 presents the average AUC for $9$ datasets with more than $10$ classes from the OpenML-CC18 benchmark suite (Bischl et al., 2021) (see Table 4 of the appendix for a list of the $9$ datasets). This represents an out-of-distribution scenario, as the pretraining phase only included datasets with up to $10$ classes. However, since our architecture is equivariant, it can handle any number of classes during inference, within memory constraints. Notably, EquiTabPFN outperforms the baselines, even with the number of classes exceeding those encountered during training.
It is worth emphasizing that performing inference on datasets with more classes than those seen during pre-training would be challenging for TabPFN and MotherNet. This limitation arises because the method would need to project vectors with dimensions exceeding those encountered during training.
6.5 Additional experiment details
The results of Tables 1 and 2 are obtained using the same evaluation protocol as in Hollmann et al. (2023) with code for training the baseline methods and selecting their hyper-parameters provided by the authors of (Müller et al., 2023) [^2] https://github.com/microsoft/ticl. Specifically, each dataset is split randomly into training and test subsets of the same size ($N=M=1000$). Predictions for each method are averaged over $5$ random splits of each dataset. The three pre-trained models we consider (EquiTabPFN, TabPFN, MotherNet) do not require any dataset specific tuning, while all other baselines are trained by allocating 60 minutes budget for hyper-parameter selection for each dataset split as done in Hollmann et al. (2023).
inline]@Michael: can you list the tids? And the number of classes considered? Are we using ensemble? No ensembles

### v2

@@ 5.1 Optimality of target equivariant networks
When presenting a new unseen dataset of covariate/target pairs $(x_{n},y_{n})_{n=1}^{N}$ to a pre-trained model, the order of the component’s target is arbitrary. In other words, given target vectors of the form $y_{n}=\left((y_{n})_{1},\dots,(y_{n})_{q}\right)$, these could as well be presented in a different order by applying a permutation $\sigma$ to the components of $y_{n}$ to obtain a permuted target $\sigma(y_{n})=\left((y_{n})_{\sigma(1)},\dots,(y_{n})_{\sigma(q)}\right)$. The transformed dataset $(x_{n},\sigma(y_{n}))_{n=1}^{N}$ is still essentially the same as the original dataset up to the permutation as we only changed the order of the target components. For instance, when the target represents a one hot encoding vector of $2$ classes: "red" or "blue", it should not matter whether we encode "red" as the first or the second class in the one-hot vector. Consequently, a pre-trained model should be able to provide consistent predictions regardless of the component’s order. More formally, the model should satisfy the following equivariance property:
@@
It is clear from Definition 5.1 that EquiTabPFN is target equivariant. Even when a model is not target equivariant by construction, it is natural to expect it to learn to be target equivariant, when trained via the objective in Equation 1 over a large class of randomly sampled datasets. To formalize this, we define the target equivariance gap, which quantifies the deviation of a function $f$ from its symmetrized counterpart ${f}^{\text{equi}}$.
Definition 5.2 (Target-equivariance gap).
The target-equivariance gap $\mathcal{E}^{\text{equi}}(f)$ of a function $f$ w.r.t. to $\mathcal{L}$ is the difference between the objective values at $f$ and its symmetrized version ${f}^{\text{equi}}$:
$\mathcal{E}^{\text{equi}}(f):=\mathcal{L}(f)-\mathcal{L}({f}^{\text{equi}}),$ (3)
where ${f}^{\text{equi}}$ is obtained by applying the following averaging operation w.r.t. to the uniform distribution $\mathbb{E}_{\sigma}$ over all permutations $\sigma$ of the target:
${f}^{\text{equi}}_{X,Y}\left(X^{\star}\right)=\mathbb{E}_{\sigma}\left[\sigma^{-1}\left(f_{X,\sigma(Y)}\left(X^{\star}\right)\right)\right].$ (4)
By construction, ${f}^{\text{equi}}$ is permutation equivariant w.r.t. the target components. Moreover it can be easily shown that a function $f$ is itself equivariant iff ${f}^{\text{equi}}=f$, so that the gap vanishes. In general, the equivariance gap can take negative values. However, we establish later that this equivariance gap must be non-negative under the following assumptions on the pointwise loss $\ell$ and the marginal distribution $p$ of the data:
@@
    Invariance and convexity of the pointwise loss. The pointwise loss $\ell$ is strictly convex in its first argument and is invariant to permutations of the components of its arguments, i.e. for any permutation $\sigma$, it holds that: $\ell(\sigma(y),\sigma(y^{\prime}))=\ell(y,y^{\prime}).$
@@
    Invariance of the data distribution. The marginal distribution of the data is invariant to permutations applied to $Y$ and $Y^{\star}$, i.e.: $p(X,\sigma(Y),X^{\star},\sigma(Y^{\star}))=p(X,Y,X^{\star},Y^{\star})$.
@@
Under Assumptions (A) and (B), the equivariance gap $\mathcal{E}^{\text{equi}}(f)$ is always non-negative and only equal to 00 when $f$ is equivariant to permutations, so that for any $f$:
$\mathcal{L}(f)=\mathcal{L}({f}^{\text{equi}})+\mathcal{E}^{\text{equi}}(f)\geq\mathcal{L}({f}^{\text{equi}}).$
Moreover, if $f^{\star}$ is a minimizer of $\mathcal{L}$ over all measurable functions, then $f^{\star}$ must be target equivariant.
@@
The key step is to express the objective as an expectation over permutations using Assumptions (A) and (B) on the data and loss: $\mathcal{L}(f)=\mathbb{E}_{p}\mathbb{E}_{\sigma}\left[\ell\left(\sigma^{-1}f_{X,\sigma(Y)}\left(X^{\star}\right),Y^{\star}\right)\right]$. The non-negativity of the equivariance gap is then established using Jensen’s inequality by convexity of the loss $\ell$ in its first argument (Assumption (A)). Now, assume by contradiction that $f^{\star}$ is not equivariant and note that ${f^{\star}}^{\text{equi}}$ is a measurable function by construction. It follows that $\mathcal{E}^{\text{equi}}(f^{\star})>0$, which directly implies that $\mathcal{L}(f^{\star})>\mathcal{L}({f^{\star}}^{\text{equi}})$, thus contradicting the optimality of $f^{\star}$. ∎
@@
Proposition 5.3 shows that minimizing the objective $\mathcal{L}$ results in a target equivariant function. Hence, using a non-equivariant model must employ some of its expressive power solely for being equivariant, which can be wasteful, as we verify empirically in Section 5.2 below.
5.2 Non-equivariance of TabPFNs models
PFNs models, as introduced in Hollmann et al. [9, 10] are not permutation equivariant in the target’s components. Consequently, they are not guaranteed to provide consistent predictions when the target components are permuted, thus affecting their robustness.
Figure 2: Prediction comparison of TabPFN, TabPFN-v2, and our model on the same datasets with three different class orderings (one per column). Models predict on a dense grid using 9 distinct training points, marked with dark crosses, each having a distinct class.
Figure 3: Equivariance error for TabPFN observed while training (top) and at inference with different number of classes and ensembles (bottom).
Predictions instabilities. To illustrate the implications of non-equivariance on the robustness of PFNs models, we consider a toy classification problem in $2$ dimensions, where $9$ training points are positioned on a $2$-dimensional regular grid, each corresponding to a different class inspired by McCarter [15]. Different pretrained models are then used to predict the classes on a regular grid of $40^{2}$ points. Figure 3 shows the classification boundaries when using the same data but with $3$ different orderings for the classes. It is clear that the ordering heavily affects the prediction even in this simple example, which strongly impacts robustness of models like TabPFN and TabPFN-v2. Note that the predictions of TabPFN-v2 are particularly noisy due to having only 9 training data points which is not an issue in itself, unlike the extreme unstability to the class ordering.
Target equivariance gap during training. In Figure 3, we analyse the equivariance error while training TabPFN. Figure 3 (left) shows the equivariance error of TabPFN in terms of percentage of violation of Equation 2, e.g. how frequently the predicted classes $f_{X,\sigma(Y)}(X^{*})$ and $\sigma\left(f_{X,Y}(X^{*})\right)$ differ. We sample 512 datasets from the prior and report standard deviation. In Figure 3 (left), the equivariance error is clear and slowly decreases during training. This non-equivariance 1) induces additional errors for the model as demonstrated in Proposition 5.3 and 2) causes the model to provide surprising predictions given that permuting the output order can change the results as seen in Figure 3.
Mitigation via costly ensembling. Hollmann et al. [9] introduce a mitigation strategy to non-equivariance by averaging several predictions using random permutations of the target dimensions. Averaging over all possible permutations gives an equivariant function as discussed in Equation 4. However, this requires making $\mathcal{O}(q!)$ calls to the model, where $q$ is the number of classes. This becomes quickly prohibitive, even for $q=10$ as considered in the original study. We illustrate how fast the model becomes equivariant when using ensembles in Figure 3 (right). While ensembling more permutation helps to make the model more equivariant, many ensembles are required, in particular, when considering more classes. In contrast, the model we propose is fully target equivariant so that predictions remain the same regardless of the class ordering as illustrated in Figure 3 (bottom).
@@ 6.1 Experimental setup
Pretraining.
We trained our model, EquiTabPFN on artificial dataset extending the public code of Müller et al. [17] and using the same procedure from Hollmann et al. [9] employed to train TabPFNv1. Crucially, it was trained on classification tasks with less than $10$ classes. The total training time took approximately 4 days on a single A100 GPU with 80GB of memory. We set our architectures hyperparameters so that they match the number of parameters of previous work, in particular the model we trained contains $\sim$25M parameters similar to the baselines we consider and we match most our hyperparameters to the same values as Hollmann et al. [9]. We refer to Appendix B for further training details and description of the hyperparameters used.
Benchmarks.
For evaluation, we consider classification tasks from the TabZilla benchmark [16]. To assess the impact of unseen class counts during evaluation (i.e., exceeding $10$), we consider two setups: one on $76$ multi-class tasks with at most $10$ classes, and another on $10$ tasks with more than $10$ classes. Details of the datasets/tasks used are provided in Tables 1 and 2. In all results, except those of Figure 5 (left), we report the results of EquiTabPFN without ensembling as it is not critical for the performance of our method.
Baselines.
We consider baselines from the TabZilla benchmark. In addition, we compare with TabPFNv2, the state-of-the-art model with open-weights released by [10] and which was shown to out-perform standard baselines, including those appearing in the TabZilla benchmark. This model was pretrained using an improved prior compared to initial version TabPFNv1. However, the code for the prior and the training is not publicly available. Hence, to allow a fair comparaison, we also include a second version, TabPFNv2, using the exact same architecture as TabPFNv2, but which we trained on the same publicly available prior and training code of [17] that we used for our model.
Figure 4: Relative improvement over KNN for datasets with less than 10 classes (left) and more than 10 classes (right). Red lines are the median metric over datasets after averaging each dataset over $10$ splits. The runtime is displayed with color on a log scale and is reported on a V100 GPU for PFNs.
6.2 Results
[…]
EquiTabPFN enables in-context-learning on datasets with unseen class counts. Figure 4 shows accuracies relative to the KNN baseline for all models on $76$ datasets with less than $10$ classes (left) and 10 datasets with more than $10$ classes (right). When considering datasets with more than $10$ classes (right), EquiTabPFN obtains the best median relative accuracy across all datasets. It strongly outperforms TabPFNv2, which performs worse than a linear model or random forests in terms of relative median. These results show that target-equivariance allows EquiTabPFN to seamlessly generalize to larger numbers of classes even though it was trained on datasets with less than $10$ classes, just like TabPFNv2. Additional evaluations in Tables 3 and 4 of Appendix C show that such improvement is consistent over various metrics: AUC, Accuracy and F1 score, for datasets with unseen class counts.
Competitive performance on datasets with class counts seen during pre-training. On the 76 datasets with less than 10 classes (Figure 4, left), EquiTabPFN performs comparably to the state-of the art method, TabPFNv2, even though it did not benefit from the improved data prior used in Hollmann et al. [10] as it is not publicly available. To assess the impact of the pre-training prior on performance, we pre-trained the same network architecture as TabPFNv2 using the same publicly available protocol used for training our method and reported the results for reference (TabPFNv2). The results show that EquiTabPFN consistently outperforms TabPFNv2 both on dataset with less than $10$ classes and those with unseen class count. These results suggest that EquiTabPFN would likely benefit from the improved training procedure and prior of TabPFNv2.
Speedup over TabPFNv2 on datasets with unseen class counts. As discussed in Section 3, PFNs models cannot natively handle an arbitrary number of classes due to its decoder architecture. In order to apply TabPFNv2 to problems with more than $10$ classes, we employ the error-correcting output codes (ECOC) strategy [4] as recommended in Hollmann et al. [10]. Such approach requires decomposing the classification problem into several smaller classification tasks with up to $10$ classes, performing predictions for each sub-task using an ensemble of TabPFNv2 models, at least one model per task, then aggregating predictions using an ECOC strategy. This incurs an extra computational cost for performing ensembling. All results reported in Figure 4 (right) are using the minimal number of ensembles to guarantee coverages of all classes. Despite this sophisticated aggregation strategy, the performance of TabPFNv2 degrades significantly, while still incurring a substantial slow-down compared to EquiTabPFN as shown by the average run-times reported by color in Figure 4 (left). Note that, for datasets with less than $10$ classes (Figure 4, left) ensembling is not required anymore by TabPFNv2 and the runtime of both methods are comparable.
EquiTabPFN achieves the best tradeoff between performance and cost on datasets with unseen class counts. To further illustrate the improved trade-off between runtime and accuracy of our method, we show in Figure 5 (left) the improvement in AUC relatively to KNN for all methods when both TabPFNv2 and EquiTabPFN are allowed to have more ensembles. Ensembling generally improves performance of PFN models, as it allows to make more robust predictions by averaging them over transformed version of the datasets, ex: by applying permutations of the labels order as discussed in Section 5.2. While ensembling helps improve performance of TabPFNv2, the improvement is marginal compared to EquiTabPFN without ensembling and comes at a considerable computational cost, as also shown in Figure 5 (right) and on the critical difference diagrams in Figures 8 and 7 of Appendix C. Therefore, amongst top performing methods, EquiTabPFN achieves the best trade-off.